OpenStudy (anonymous):

simply (v + 2) (v-5)

OpenStudy (anonymous):

Use FOIL: multiply v by v \[v \times v=v ^{2}\] multiply v by -5 \[v \times -5=-5v\] multiply 2 by v \[2 \times v=2v\] multiply 2 by -5 \[2 \times -5=-10\] add terms \[v ^{2}-5v+2v-10\] simplify \[v ^{2}-3v-10\]